# Federated Learning for Surgical Vision in Appendicitis Classification: Results of the FedSurg EndoVis 2024 Challenge

### v1

@@ 4 Discussion
4.1 Task Wise performance
The FedSurg Challenge establishes the first benchmark for evaluating Federated Learning (FL) strategies in the surgical video classification of appendicitis grades, utilizing a heterogeneous, multi-institutional dataset. Three complete submissions were received and evaluated on a preliminary subset of the Appendix300 dataset.
Task 1, which focused on generalization to an unseen center, immediately revealed the limited generalization abilities of current FL models. This difficulty stems directly from significant data heterogeneity across the participating institutions. For instance, images originated from a wide array of surgical video systems, including Arthrex Synergy UHD4, Storz image1 s, B. Braun Aesculap EinsteinVision, and Richard Wolf Endocam 4k, leading to variations in lighting and resolution [21]. For example, Center 1 provided 4k images, while Center 4 provided cropped images in Full HD resolution. In addition to this feature skew, a label distribution skew existed between the different centers (Figure 2). In this challenging environment, models capable of temporal modeling, like the ViViT-based submission, achieved the strongest but still weak results. A slight tendency also emerged showing that approaches focused on the loss landscape, like FedSAM, performed slightly better than a FedMedian aggregation of a triplet margin loss and cosine similarity model. This aligns with the conclusions of Foret et al. [39] and Caldarola et al. [38] that the quest for flat minima, which is incorporated into FedSAM, can lead to better generalization ability. This difficulty was not unique to the challenge submissions; a benchmark study on the Appendix300 dataset [22] using a leave-one-out cross-validation design similarly confirmed the profound challenge of inter-center generalization.
Task 2, which involved center-specific adaptation after fine-tuning, led to only marginally better results across all teams. This marginal improvement, however, came at the cost of rank instability, highlighting that even minor gains in local adaptation can compromise the robustness of a model’s comparative performance. The limited improvement can be attributed to several factors inherent to the federated setup. Due to the difficult knowledge of the complete data distribution and amounts, hyperparameter tuning was challenging, which could lead to local models overfitting to their local training data and, therefore, still failing to achieve high accuracy. Furthermore, the knowledge from other centers within the global model can negatively influence client performance. At each training round, the client initializes its weights from the global model, which represents an aggregated compromise of the knowledge learned across all participating centers. The poor local performance could stem from conflicting updates that pull the model away from what is optimal for its own center’s data.
Ultimately, these findings illustrate the central challenge of balancing generalization and adaptation, alongside the overarching difficulty of data heterogeneity in federated setups.
4.2 Stability assessment
The bootstrapped rank frequency analysis yielded distinct outcomes for Task 1 and Task 2.
For Task 1, the analysis indicated stable leader board rankings, with the predictions from submitted models being statistically distinct. A central finding was the critical role of temporal information; the highest-ranking submission utilized a video-based model, whereas models predicated on static frames exhibited the poorest performance. This aligns with the findings from the Appendix300 benchmarking paper [22], where the model trained only on middle frames performed poorly in comparison to the models trained on video models with a temporal horizon. Furthermore, using a higher frame rate up to 1 frame per second (fps) resulted into better results.
In stark contrast, the rankings for Task 2 demonstrated pronounced instability. No submission maintained its rank on the EC metric in more than 50% of the bootstrap replicates, indicating that the leader board outcomes were not robust. This instability suggests that leader boards for federated adaptation tasks should be interpreted with extreme caution, as top-ranking methods may not be reliably superior but merely fortunate in a specific data configuration.
4.3 Systemic Limitations and Data-Related Challenges in Federated Surgical AI
Several systemic issues emerged. FL does not resolve underlying dataset problems such as class imbalance and can amplify heterogeneity effects. Rare inflammation stages were underrepresented, leading to poor classification across all teams. Personalized losses and fine-tuning improved performance within centers (e.g., Team Camma) but compromised generalization to unseen centers, underscoring a fundamental trade-off in FL. This mirrors our findings from the primary tasks, where the unstable gains in center-specific adaptation (Task 2) failed to translate into the robust, generalizable models needed for unseen centers (Task 1). Similarly, design choices impacted outcomes: lightweight frozen foundation models provided stability across centers, whereas more adaptive approaches (e.g., Siamese networks with triplet loss) improved local performance but struggled with generalization under distribution shifts.
Furthermore, hyperparameter tuning proved difficult without access to the true data distribution, reflecting real-world FL constraints. The unknown class distribution contributed to lower-than-expected performance. Moreover, human experts showed only moderate agreement on inflammation grading ($\kappa_{weighted}=0.62$ [21]), underscoring the inherent ambiguity of the classification task.
4.4 Future directions
Future work should address identified limitations. Imbalance-aware strategies, such as reweighting, resampling, or federated focal loss, are needed to improve performance on clinically critical minority classes. Context-aware preprocessing and targeted frame sampling consistently outperformed uniform strategies and should be further explored. Methodological advances such as federated domain adaptation, personalized aggregation, and uncertainty-aware inference may improve robustness and safety. Finally, self-supervised pretraining on large-scale surgical video datasets could yield more transferable representations without compromising data privacy.
4.5 Challenge insights
Despite 24 registered teams, only three complete submissions were ultimately received. This discrepancy reflects both the niche character of FL within surgical AI and the complexity of the challenge setup. Unlike conventional challenges, FedSurg required fully encapsulated end-to-end submissions in Docker, with limited access to data and restricted debugging capabilities. These hurdles highlight practical barriers to broader participation and point to the technical demands of real-world FL applications.
@@ 5 Conclusion
The FedSurg Challenge provides the first benchmark for FL in surgical video classification, highlighting both opportunities and current limitations. The findings demonstrate that while spatiotemporal modeling and preprocessing choices can support generalization and adaptation, challenges persist in handling data imbalance, hyperparameter tuning, and stability of results. By exposing these trade-offs, FedSurg establishes an important reference point for the surgical AI community and guides future work toward robust, adaptive, and clinically relevant FL methods.
@@ Acknowledgments
During the preparation of this work, the author(s) used ChatGPT, Gemini 2.5 Pro and DeepL Write in order to improve the readability and language of the manuscript. After using these tools/services, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Appendix A Dataset and Annotation Protocol
The challenge dataset is a preliminary subset of the publicly available Appendix300 dataset [21]. A CSV file detailing the samples used in the challenge is available here  [^2] Available upon acceptance or on request., along with two videos that were excluded from the final dataset but used in the challenge for reproducibility. The dataset is available for non-commercial use under a CC BY license and requires citation of the Appendix300 publication. The annotation protocol is available in [21]. Usage of the data for individual publications was prohibited before the release of this study and the Appendix300 dataset.
@@ Appendix C Challenge Organization
This work was jointly organized by four parties. Fiona R. Kolbinger, from the Department of Visceral, Thoracic and Vascular Surgery, Faculty of Medicine and University Hospital Carl Gustav Carus of the TUD Dresden University of Technology and the Weldon School of Biomedical Engineering, Purdue University coordinated the challenge data collection from four hospitals in Germany. Oliver S. Lestner and Jakob N. Kather, also from the Else Kröner Fresenius Center for Digital Health, supported the data aggregation from the technical point of view and obtained the source data from these centers. Max Kirchner, Alexander C. Jenke, Sebastian Bodenstedt, and Stefanie Speidel, from the Department of Translational Surgical Oncology (TSO) of the National Center for Tumor Diseases (NCT) Dresden, supported data aggregation and carried out the challenge organization and the technical implementation, including preprocessing, data provision, participant registration and administration, submission handling, evaluation, and results presentation. The NearData Horizon Europe program provided €500 in prize money, equally distributed across both tasks. If a single team achieved the top score in both tasks, they were eligible to receive both awards.
@@ Appendix G Additional Bootstrapping and Wilcoxon Signed-Rank Test Results
Appendix H Credit Authorship Contribution Statement

### v2

@@ 4 Discussion
The FedSurg challenge should be understood as a proof-of-concept benchmark: the appendectomy task was selected for its methodological tractability rather than for its immediate clinical impact. It offers a standardized surgical workflow, multi-center data availability, and an established grading system. The following discussion focuses on methodological insights transferable to future FL applications on surgical video, structured around two central questions: first, what architectural and algorithmic choices best support global generalization to unseen centers; and second, whether and how local adaptation can be reliably achieved without sacrificing the globally learned representation.
4.1 The Role of Temporal Modeling
A key finding of this challenge is that temporal modeling strategy is the dominant performance factor, outweighing the choice of aggregation scheme or communication frequency. Frame-level approaches, whether operating through majority voting (Team Elbflorenz) or embedding-based prototype comparison (Team Camma), are inherently limited by the high intra-video visual variance present in laparoscopic recordings. Differences in laparoscopic hardware, lighting conditions, and patient anatomy render individual frames unreliable predictors of patient-level inflammation grade. The metric learning approach of Team Camma illustrates this most clearly: while it achieved competitive performance at centers whose visual characteristics were represented in training, it failed to map embeddings of the unseen center to established training prototypes, revealing the brittleness of a feature space built on individual frames.
Video-level temporal models, by contrast, capture the progression of visual features across the surgical sequence, providing contextual information no single frame can supply. The ViViT-based submission of Team Santhi, processing 32 frames jointly, consistently achieved the strongest generalization to the unseen center in Task 1, outperforming both frame-level approaches despite using fewer communication resources. This finding is consistent with prior work on decentralized surgical video analysis using the same dataset [42], in which temporal architectures similarly outperformed frame-level methods across patient-level prediction tasks.
With respect to the generalization vs. adaptation trade-off, temporal modeling also offered a partial advantage in Task 2: Team Santhi’s frozen ViViT backbone preserved global spatiotemporal representations during local fine-tuning, limiting the degree to which center-specific updates could corrupt the globally learned feature space. This suggests that architecture choice is not only a generalization strategy but also a safeguard against the instability of local adaptation. This point is revisited in Section 4.4.
4.2 Dataset Scale and Decentralization as Compounding Factors
The generally low absolute scores across all submissions raise an important question: do these results reflect fundamental limitations of FL, or primarily the difficulty of the task itself? The centralized baseline provides a direct answer. Even when pooling all multi-institutional data onto a single server, thereby removing all federated constraints, the centralized model achieved only 26.31% F1-score on the unseen center. This establishes that the underlying task of patient-level appendicitis grading from surgical video is intrinsically difficult, regardless of the learning paradigm employed, due to the high visual variance across centers and the inherent ambiguity of fine-grained clinical grading.
Decentralization acts as an additional, separable bottleneck. The Swarm Learning baseline, which retains the same architecture as the centralized model but distributes training across centers, achieved 19.00% F1 in Task 1. This measurable degradation is attributable specifically to decentralized optimization across non-IID data nodes, rather than to any architectural deficiency. This gap quantifies what may be termed a decentralization penalty: the cost in predictive performance of preserving data privacy through distributed training. Together, the two baselines confirm that observed performance reflects a compounding of two distinct factors, and that algorithmic advances in FL aggregation must overcome both.
An additional consideration is the scale of the dataset. The challenge utilized 153 training videos, which constitute a preliminary subset of the 330-video Appendix300 dataset [24], distributed across four centers. At this data regime, FL fragments an already limited training pool, leaving individual nodes with insufficient samples to learn robust local representations. Prior work using the full dataset [42] demonstrated improved performance under decentralized training, suggesting that the current performance ceiling is partly a function of data scarcity rather than an inherent ceiling of the FL paradigm. Scaling the dataset, in terms of both total videos and number of contributing centers, is therefore a necessary condition for a more definitive evaluation of FL on surgical video.
4.3 Data Heterogeneity and the Limits of Local Adaptation
The challenge results highlight that non-IID data remains the central unsolved problem in federated surgical video analysis. Distribution shift across centers originates from two compounding sources: hardware-induced feature skew, including differences in laparoscopic systems, resolution, lighting, and cropping, and patient population heterogeneity, including differences in age distribution, disease severity prevalence, and surgical practice between academic and community centers. Current aggregation strategies, including FedAvg and FedSAM, partially mitigate but do not resolve these shifts.
These effects are further amplified by inter-center data quantity skew. Center 3 contributed 80 of the 153 training videos, causing global models to disproportionately reflect Center 3’s data distribution. The bootstrapped rank stability analysis illustrates the consequence: Task 1 rankings were stable, confirming that temporally-informed models were statistically distinguishable; Task 2 rankings showed pronounced instability, with no submission retaining its rank on the EC metric in more than 50% of bootstrap replicates. Local fine-tuning at smaller centers triggered reversion to trivial classifiers, predominantly predicting a single class as a consequence of catastrophic forgetting on heavily skewed local data, inflating nominal F1-scores without reflecting genuine adaptation. As discussed in Section 2.4, the EC metric was more robust to this artifact and should be considered the primary indicator of adaptation quality in imbalanced ordinal settings.
4.4 Toward Personalized Federated Learning
The results of both tasks point toward the same conclusion: a single globally aggregated model is a compromise that satisfies no center optimally. However, naive local fine-tuning, as evaluated in Task 2, does not resolve this, as unconstrained adaptation on skewed local data leads to catastrophic forgetting rather than genuine personalization. The path forward lies in structured personalized FL: approaches that separate a shared global representation from center-specific adaptation components, and train both with objectives that explicitly account for local data imbalance.
For Task 1 generalization, the FedSurg results support the use of a frozen pretrained temporal backbone as the globally federated component. By fixing backbone weights across federation rounds, center-specific visual variance cannot propagate into the shared representation, preserving its generalizability to unseen centers. Only lightweight task-specific heads need to be aggregated globally, substantially reducing both communication cost and the risk of conflicting updates between heterogeneous nodes.
For Task 2 adaptation, parameter-efficient fine-tuning (PEFT) methods offer a principled alternative to full local fine-tuning. Methods such as LoRA or adapter layers constrain the parameter space available for local updates, limiting the degree to which a small, imbalanced local dataset can destabilize the global representation, while still enabling meaningful center-specific adjustment. This approach is particularly relevant given the data regime of this challenge, where local node datasets ranged from 33 to 80 videos.
Looking further ahead, surgical video foundation models represent a promising direction for initializing the shared backbone [48, 43]. However, currently available surgical foundation models are trained on limited procedure-specific data, and general-purpose video foundation models produce feature spaces too broad for the fine-grained visual distinctions required in intraoperative grading tasks. Recent work combining FL with endoscopic foundation model pretraining [22] suggests a more principled path: federating the pretraining process itself to build privacy-preserving, domain-specific representations from the outset. The integration of such approaches with PEFT-based local adaptation remains a hypothesis for future evaluation rather than a validated recommendation.
The exploratory PEFT baselines included in this work provide initial empirical grounding for these recommendations. The EndoViT+LoRA variant, which applies LoRA to a frozen frame-level backbone, collapsed to majority-class prediction (Task 1 F1: 9.24%), confirming that parameter-efficient adaptation cannot compensate for the absence of temporal modeling in this data regime. The ViViT+LoRA variant, applying LoRA to a frozen temporal backbone, achieved 15.93% F1 on Task 1 and 29.57% F1 at Center 3 in Task 2, approaching the centralized baseline (32.67%) at the largest local node. Performance at smaller centers remained limited (Center 1: 12.22%, Center 2: 22.62%). This is consistent with the quantity-skew hypothesis: PEFT-based local adaptation is most effective where sufficient local samples are available to support stable fine-tuning.
4.5 Implications for Task Design
Beyond algorithmic choices, the FedSurg results suggest that the design of the clinical prediction task itself should inform the choice of FL strategy. This dimension has received limited attention in the literature. The consequences of prediction errors differ fundamentally between tasks, and these differences have direct implications for how generalization and adaptation should be balanced.
Consider two contrasting task types. In high-sensitivity settings, for example the detection of malignant lesions in colonoscopy where false negatives carry direct patient harm, global generalization to unseen centers is critical. FL strategies that prioritize robustness across non-IID distributions, such as flat-minima optimization approaches like FedSAM, are preferable over those that optimize local performance at the cost of cross-center generalizability. The fine-grained 6-class inflammation grading evaluated in this challenge approaches the opposite end of the spectrum: overcalling severity could influence clinical decision-making, placing greater weight on specificity. Here, the instability of local adaptation observed in Task 2 is particularly concerning. Classifier collapse, the reversion to a trivial classifier predominantly predicting a single class as a consequence of catastrophic forgetting during fine-tuning on imbalanced local data, is not only a performance failure but a risk in high-specificity grading contexts.
These observations lead to a practical hypothesis for future work: FL strategy selection should be guided not only by dataset characteristics, but by an explicit analysis of the clinical error cost structure of the target task. High-sensitivity tasks call for globally robust aggregation; high-specificity tasks require adaptation strategies that prevent classifier collapse, such as the PEFT-based approaches discussed in Section 4.4. Systematic evaluation of this hypothesis across diverse surgical AI tasks remains an important open problem.
4.6 Limitations and Future Directions
Several limitations of this challenge should be acknowledged. Despite 24 registered teams, only three complete submissions were received, reflecting both the niche character of FL within surgical AI and the high technical barriers to participation. The dataset, as a preliminary subset of Appendix300 [24], was insufficient in scale for robust federated training, and the simulated FL setup, where all data resided centrally during evaluation, does not capture the full complexity of real-world distributed deployments, including network latency, node availability, and hardware heterogeneity. Future work should prioritize scaling the dataset across a broader network of institutions and transitioning from simulated to genuine multi-institutional FL. On the algorithmic side, imbalance-aware objectives, such as federated focal loss or client-weighted aggregation, and personalized aggregation methods that account for quantity skew are necessary next steps. Self-supervised pretraining on large-scale surgical video collections, ideally in a federated manner [22], offers a promising route toward more transferable representations without compromising data privacy.
@@ 5 Conclusion
The FedSurg Challenge establishes the first benchmark for FL in surgical video classification, evaluated here as a proof-of-concept using a multi-center dataset of laparoscopic appendectomies. The results reveal that temporal modeling is the dominant architectural factor for cross-center generalization, while decentralized training introduces a measurable performance penalty on top of the intrinsic difficulty of the task. Local fine-tuning without imbalance-aware objectives leads to classifier collapse rather than genuine personalization, underscoring the need for structured personalized FL with parameter-efficient adaptation. By characterizing these failure modes through rigorous statistical analysis, this work provides a methodological reference point for the design of robust, privacy-preserving AI systems for surgical video analysis and motivates future benchmarking on tasks with broader clinical scope.
@@ Acknowledgments
During the preparation of this work, the author(s) used ChatGPT, Gemini 2.5 Pro, Claude Code, and DeepL Write in order to improve the readability and language of the manuscript. After using these tools/services, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Appendix A Dataset and Annotation Protocol
The challenge dataset is a preliminary subset of the publicly available Appendix300 dataset [24]. A CSV file detailing the samples used in the challenge is available here  [^2] Available at https://nextcloud.tso.ukdd.de/s/XppWkRnWpLGwj5F Password for Review: FedSurg24_data!, along with two videos that were excluded from the final dataset but used in the challenge for reproducibility. The dataset is available for non-commercial use under a CC BY license and requires citation of the Appendix300 publication. The annotation protocol is available in [24]. Usage of the data for individual publications was prohibited before the release of this study and the Appendix300 dataset.
@@ Appendix C Challenge Organization
This work was jointly organized by four parties. Fiona R. Kolbinger, from the Department of Visceral, Thoracic and Vascular Surgery, Faculty of Medicine and University Hospital Carl Gustav Carus of the TUD Dresden University of Technology and the Weldon School of Biomedical Engineering, Purdue University coordinated the challenge data collection from four hospitals in Germany. Oliver L. Saldanha and Jakob N. Kather, also from the Else Kröner Fresenius Center for Digital Health, supported the data aggregation from the technical point of view and obtained the source data from these centers. Max Kirchner, Alexander C. Jenke, Sebastian Bodenstedt, and Stefanie Speidel, from the Department of Translational Surgical Oncology (TSO) of the National Center for Tumor Diseases (NCT) Dresden, supported data aggregation and carried out the challenge organization and the technical implementation, including preprocessing, data provision, participant registration and administration, submission handling, evaluation, and results presentation. The NearData Horizon Europe program provided €500 in prize money, equally distributed across both tasks. If a single team achieved the top score in both tasks, they were eligible to receive both awards.
@@ Appendix G Additional Bootstrapping and Wilcoxon Signed-Rank Test Results
Appendix H Further Analysis Results with Baselines
Table 11: Comprehensive ranking summary for Task 1, Task 2, and overall performance including contextual baselines. Lower rank indicates better performance.
Team Task 1 Task 2 Overall Final
EC F1 Avg EC F1 Avg Avg Rank
Camma 5 5 5 5 2 3 4.00 3
Elbflorenz 4 4 4 4 5 4 4.00 3
Santhi 2 2 2 3 4 3 2.50 2
Cen 1 1 1 1 1 1 1.00 1
SWARM 3 3 3 2 3 2 2.50 2
H.0.1 Detailed Bootstrap Analysis Tables
Task 1: Generalization (F1-score)
Tables 1214 report rank frequency, pairwise win probabilities, and Wilcoxon p-values for Task 1 F1-score. Rankings are stable: the centralized baseline holds rank 1 in 67.0% of bootstrap iterations and all pairwise comparisons are statistically significant ($p<0.001$).
Table 12: Task 1 F1-score: Rank frequency distribution.
Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
Camma 0.0000 0.0000 0.0007 0.2486 0.7507
Elbflorenz 0.0002 0.0022 0.0111 0.7385 0.2480
Santhi 0.2961 0.4736 0.2233 0.0070 0.0000
Cen 0.6700 0.2917 0.0377 0.0005 0.0001
SWARM 0.0350 0.2319 0.7265 0.0065 0.0001
Table 13: Task 1 F1-score: Pairwise win probabilities $P(\text{team}_{i}>\text{team}_{j})$.
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.2481 0.0000 0.0001 0.0007
Elbflorenz 0.7509 0.0077 0.0011 0.0074
Santhi 1.0000 0.9923 0.3024 0.7637
Cen 0.9999 0.9989 0.6974 0.9330
SWARM 0.9992 0.9926 0.2361 0.0654
Table 14: Task 1 F1-score: Wilcoxon p-values. All comparisons $p<0.001$.
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.000 0.000 0.000 0.000
Elbflorenz 0.000 0.000 0.000 0.000
Santhi 0.000 0.000 0.000 0.000
Cen 0.000 0.000 0.000 0.000
SWARM 0.000 0.000 0.000 0.000
Task 1: Generalization (Expected Cost)
Tables 1517 report the equivalent analysis for the EC metric. Rankings are similarly stable, with Camma retaining the last rank in 100% of iterations and all pairwise differences statistically significant.
Table 15: Task 1 EC: Rank frequency distribution.
Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
Camma 0.0000 0.0000 0.0000 0.0000 1.0000
Elbflorenz 0.0000 0.0005 0.0035 0.9960 0.0000
Santhi 0.2576 0.3401 0.4003 0.0020 0.0000
Cen 0.7384 0.2480 0.0136 0.0000 0.0000
SWARM 0.1021 0.3715 0.5254 0.0010 0.0000
Table 16: Task 1 EC: Pairwise win probabilities.
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.0000 0.0000 0.0000 0.0000
Elbflorenz 1.0000 0.0021 0.0001 0.0013
Santhi 1.0000 0.9973 0.2048 0.5382
Cen 1.0000 0.9999 0.7355 0.8858
SWARM 1.0000 0.9983 0.4091 0.0703
Table 17: Task 1 EC: Wilcoxon p-values. All highly significant ($p<0.001$).
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.000 0.000 0.000 0.000
Elbflorenz 0.000 0.000 0.000 0.000
Santhi 0.000 0.000 0.000 $2.89\times 10^{-65}$
Cen 0.000 0.000 0.000 0.000
SWARM 0.000 0.000 $2.89\times 10^{-65}$ 0.000
Task 2: Adaptation (F1-score)
Tables 1820 show markedly higher rank volatility compared to Task 1. No team retains a single rank in more than 66.4% of iterations, and the pairwise win probabilities between Camma, Santhi, and SL indicate substantial overlap, reflecting the instability of local adaptation across centers.
Table 18: Task 2 F1-score: Rank frequency distribution.
Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
Camma 0.1794 0.3490 0.2499 0.1517 0.0700
Elbflorenz 0.0078 0.0517 0.1407 0.3063 0.4935
Santhi 0.0153 0.0992 0.2188 0.3254 0.3413
Cen 0.6641 0.1999 0.0866 0.0367 0.0127
SWARM 0.1335 0.3001 0.3041 0.1798 0.0825
Table 19: Task 2 F1-score: Pairwise win probabilities.
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.8507 0.7786 0.2326 0.5540
Elbflorenz 0.1493 0.4091 0.0408 0.1748
Santhi 0.2212 0.5909 0.0648 0.2447
Cen 0.7674 0.9592 0.9352 0.8042
SWARM 0.4460 0.8252 0.7553 0.1958
Table 20: Task 2 F1-score: Wilcoxon p-values.
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.000 0.000 0.000 $9.54\times 10^{-35}$
Elbflorenz 0.000 0.000 0.000 0.000
Santhi 0.000 $1.49\times 10^{-101}$ 0.000 0.000
Cen 0.000 0.000 0.000 0.000
SWARM $9.54\times 10^{-35}$ 0.000 0.000 0.000
Task 2: Adaptation (Expected Cost)
Tables 2123 report Task 2 EC bootstrap results, where rank instability is most pronounced. Rank frequencies are close to uniform for most teams, confirming that EC-based adaptation rankings are not reliably reproducible under resampling and should be interpreted with caution.
Table 21: Task 2 EC: Rank frequency distribution.
Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
Camma 0.1367 0.1533 0.1737 0.2169 0.3194
Elbflorenz 0.1262 0.1773 0.2202 0.2377 0.2386
Santhi 0.2035 0.2317 0.2299 0.1956 0.1393
Cen 0.3043 0.2228 0.1840 0.1628 0.1261
SWARM 0.2323 0.2154 0.1933 0.1849 0.1741
Table 22: Task 2 EC: Pairwise win probabilities.
Team Camma Elbflorenz Santhi Cen SWARM
Camma 0.4637 0.3791 0.3333 0.3872
Elbflorenz 0.5336 0.4015 0.3571 0.4156
Santhi 0.6166 0.5954 0.4415 0.5026
Cen 0.6665 0.6422 0.5578 0.5477
SWARM 0.6123 0.5839 0.4971 0.4517
Table 23: Task 2 EC: Wilcoxon p-values. Note: Santhi vs. SWARM $p=0.746$ (not significant).
Team Camma Elbflorenz Santhi Cen SWARM
Camma $1.30\times 10^{-21}$ $2.72\times 10^{-189}$ 0.000 $1.21\times 10^{-171}$
Elbflorenz $1.30\times 10^{-21}$ $3.81\times 10^{-72}$ 0.000 $4.71\times 10^{-98}$
Santhi $2.72\times 10^{-189}$ $3.81\times 10^{-72}$ 0.000 0.746
Cen 0.000 0.000 0.000 $2.86\times 10^{-38}$
SWARM $1.21\times 10^{-171}$ $4.71\times 10^{-98}$ 0.746 $2.86\times 10^{-38}$
Appendix I Credit Authorship Contribution Statement